Find the slope over the interval [0,4]
Y=(X^3)-2x

Answer by Alan3354(69443)

Find the slope over the interval [0,4]
Y=(X^3)-2x
--------
f(x) = x^3 - 2x
---
f(4) = 64 - 8 = 56
f(0) = 0
---
(56 - 0)/4 = 14, the slope of the line between the 2 points.
----
The slope at any point is 3x^2 - 2